# Empowering Scientific Workflows with Federated Agents

### v2

@@ Abstract.
Agentic systems, in which diverse agents cooperate to tackle challenging problems, are exploding in popularity in the AI community. However, the agentic frameworks used to build these systems have not previously enabled use with research cyberinfrastructure. Here we introduce Academy, a modular and extensible middleware designed to deploy autonomous agents across the federated research ecosystem, including HPC systems, experimental facilities, and data repositories. To meet the demands of scientific computing, Academy supports asynchronous execution, heterogeneous resources, high-throughput data flows, and dynamic resource availability. It provides abstractions for expressing stateful agents, managing inter-agent coordination, and integrating computation with experimental control. We present microbenchmark results that demonstrate high performance and scalability in HPC environments. To demonstrate the breadth of applications that can be supported by agentic workflow designs, we also present case studies in materials discovery, decentralized learning, and information extraction in which agents are deployed across diverse HPC systems.
Computational Workflows, Distributed Computing, Federated Computing, Multi-Agent Systems, Open-Source Software
1. Introduction
@@
The desire to automate scientific processes has led to advancements in many fields, from artificial intelligence (AI) (Zvyagin et al., 2023) and computational workflows (Deelman et al., 2009) to research data management (Allen et al., 2012) and self-driving laboratories (SDL) (Abolhasani and Kumacheva, 2023), but humans typically remain responsible for core aspects of the iterative research cycle, including hypothesis generation, experimental design, code development, and data analysis. Often, the human-in-the-loop is the rate-limiting step in discovery. This friction increases as the scale and ambition of computational science endeavors grow and leads to inefficient use of research cyberinfrastructure—the federated ecosystem of experimental and observational facilities, data repositories, and high-performance computing (HPC) systems (Miller et al., 2023).
Figure 1. Cooperative agents, spanning federated research infrastructure (experimental facilities, computational systems, data storage), can enable agentic workflows that autonomously steer discovery.
[…]
[Agents deployed across research infrastructure, including experimental facilities, computational systems, and data storage, can facilitate agentic workflows that autonomously engage research infrastructure.]
@@
Intelligent agents, either as an individual system or composing larger multi-agent systems (MAS), rather than humans, can be the driving entities of discovery. Agents are independent, persistent, stateful, and cooperative—working together to achieve a predefined goal with only intermittent human oversight. The contemporaneous explosion of interest in multi-agent systems is largely a consequence of advancements in reasoning capabilities of the large language models (LLMs) often used to back AI agents. Expressing components of scientific applications as agents—programs that can perform tasks independently or semi-autonomously on behalf of a client or another agent—is powerful. An agent manages its own local state and exposes a well-defined behavior. Agents can perform human roles in iterative scientific processes (Wang et al., 2023) or encapsulate research cyberinfrastructure (e.g., computational resources and procedures, experimental instruments, data repositories) (Gao et al., 2024).
Significant progress has been made towards developing AI agents that can act on behalf of humans for such tasks as literature synthesis (Lála et al., 2023), hypothesis generation (Google Research, 2025), and data analysis and publication (Ifargan et al., 2025). However, existing agent frameworks (e.g., AutoGen (Wu et al., 2023)) are not ready to build and deploy agents that employ federated research cyberinfrastructure. New middleware is needed to enable agentic workflows that seamlessly integrate experiment, observation, theory, simulation, AI, analysis, and more, as in Figure 1.
Frameworks for building agentic workflows are limited in scope and generally target conversational, cloud-native applications (e.g., LLM-based AI chatbots) (LangChain, 2024; Wu et al., 2023; OpenAI, 2024). The federated nature of research infrastructure poses unique challenges: distributed resources have diverse access protocols, interactions between computational and experimental entities are asynchronous, and the dynamic availability of resources requires fault-tolerant and adaptive systems. Existing frameworks fail to address these intricacies. They lack abstractions and mechanisms tailored to support autonomous multi-agent workflows that integrate computation, data management, and experimental control, which leads to brittle, ad hoc integrations that are ill-suited for the demands of modern science. Moreover, the inherent complexity of such workflows is compounded by the need to balance efficiency with scientific rigor, especially in applications involving real-time decision-making, iterative exploration, and multi-agent coordination.
These challenges are often orthogonal and span many levels of abstraction, but achieving this vision where intelligent agents serve as driving entities in scientific discovery requires a paradigm shift in how workflows are designed, orchestrated, and executed. We introduce a novel framework for building agentic workflows, emphasizing modularity, statefulness, and interoperability across the diverse research infrastructure. Specifically, this work contributes:
@@
  • Academy, a novel, modular, and extensible middleware for expressing agentic workflows and deploying multi-agent systems across federated resources. Academy addresses unique challenges in scientific applications, such as high data volumes, variable resource availability, and the heterogeneous nature of experimental and computational systems (Section 3).
@@
  • Performance analysis of Academy in diverse scenarios yielding insights into the scalability and practical considerations of deploying agentic workflows (Section 4).
@@
  • Case studies demonstrating the utility of agentic workflow design and highlighting improvements in automation, resource utilization, and discovery acceleration (Section 5).
@@ 2. Background
Agents encompass a rapidly expanding front for AI research, yet agent paradigms can address a breadth of challenges across the computational sciences. We begin with a definition of an agent—inspired by prior work—that is sufficiently generic to encompass the various semantic uses of the term. Then, we enumerate common high-level classes of agents and formalize agentic workflows, both of which we aim to support in the design of Academy in Section 3.
@@
An agent is a program that can perform actions independently or semi-autonomously on behalf of a client or another agent. This definition is imprecise but presents a powerful conceptual model for distributed computing. The agent concept originates from the actor model, a concurrent computing paradigm in which actors encapsulate a local state and communicate through message passing (Hewitt et al., 1973). Agents extend the actor model with the notion of agency—the ability of the agent to engage independently with its environment.
An agent $a$ is defined by its behavior $B$ and local state $S$. The behavior of an agent encompasses a set of actions $x\in X$ (procedures that the agent can perform), and a set of control loops $c\in C$ that define the autonomous behavior (Goodwin, 1995). Agents are often long-running, but may also be ephemeral—created to complete a specific task and then exiting. Clients and agents can request another agent to perform an action on their behalf through message passing. An action can be atomic or composite, invoking other actions on the same or remote agents. An agent with actions but no control loops (i.e., $|X|>0$ and $|C|=0$) reduces to an actor.
Agents come in many flavors. Intelligent (deliberative) agents are goal-oriented and reason about what actions to take using internal models and external perception (Wooldridge and Jennings, 1995). AI agents, a subset of intelligent agents, use AI methods to make decisions or perform actions. In contrast, reactive (observer) agents simply perceive their external environment and react to changes (Nwana, 1996). Service agents provide predefined services in response to action requests and come in many forms: resource agents manage and grant access to resources, such as compute or storage, and embodied agents can act in the world, such as through physical actions when paired with a robot body. Learning agents adapt their behavior over time to improve performance, often through reinforcement learning (Mnih et al., 2015). Composite agents exhibit two or more of these behaviors. For example, deliberative learning agents improve their reasoning or planning capabilities over time, and reactive service agents perform services in response to environmental changes. A multi-agent system can enable more complex behaviors than monolithic programs (Panait and Luke, 2005), which can lead to powerful emergent behavior (Cucker and Smale, 2007).
An agentic workflow can be formalized as a graph of actions—rather than tasks, as in typical DAG-based workflows—where agents request and perform actions on behalf of one another, enabling dynamic coordination and the collective pursuit of complex, distributed objectives. Let the environment $E$ represent the external state space, influencing and influenced by the actions of agents and other entities. Agents in the environment are represented by a deployment $d(A,R):A\rightarrow R$ of agents $a\in A$ on to resources $r\in R$. Each agent implements a behavior, and an agent that knows the behavior of a peer agent can request the peer to perform an action through message passing. Thus, there exists a directed graph representing the peer relationships between the agents; an edge $e=(a_{i},a_{j})$ in this graph implies that $a_{i}$ knows of $a_{j}$ and can request actions from $a_{j}$. Sink nodes in the graph represent agents that only perform atomic actions, whereas source nodes may represent deliberative or reactive agents that trigger actions on other agents. A cut vertex (articulation point) in the graph can represent an agent that serves as an interface or gateway to another, possibly more complex, multi-agent system.
The deployment of agents can execute workflows. An agentic workflow $W$ is modeled as a directed graph where nodes are a tuple $(x,a)$ of an action to perform and the agent performing the action, and edges representing the source agent and action that triggered the subsequent action. A workflow is typically implicitly encoded within the agent behaviors of a multi-agent system. I.e., the graph $W$ is not explicitly materialized and agents do not need to know $W$ in order to execute. An agent only needs to be concerned with its local view of executing requested actions and requesting actions from peers. Thus, workflows may often be highly dynamic as agent behaviors react to changing states.
3. Academy Design
Designing a middleware that can express the diverse demands of scientific applications and leverage federated research infrastructure is challenging. In the design of Academy, we aim to address the following high-level challenges: How to represent, in code, the declaration of and interaction between agents? How to deploy agents across federated infrastructure? How to achieve performance across heterogeneous systems, networks, and storage? Thus, we begin by outlining key requirements, before we introduce the high-level architecture and detailed implementation choices. The name Academy alludes to societies of artists and scholars that, while independent, collaborate and share similar goals.
3.1. Academy Requirements
Writing scientific applications as agentic workflows, rather than using traditional workflow models, can require a considerable shift in conceptual thinking. To reduce developer friction, our design emphasizes familiarity—using well-known programming patterns—and simplicity—providing a small set of primitives and inviting users to invent new patterns and techniques. With these principles in mind, we define requirements in four areas:
  • Representation: Agent behavior must be expressed in code, supporting control loops, actions that can be performed, and local state. Multiple agents may be instantiated with the same behavior. Agents should not share state.
  • Interaction: Agents and clients must be able to interact. They must be able to address a specific agent, perform one or more actions, modify local state, or create and terminate agents.
  • Communication: Agents and clients communicate asynchronously and are temporally decoupled (e.g., a message sent to an offline agent should be read when that agent is next online). Agents may be deployed in diverse environments with heterogeneous network environments (e.g., asymmetric networks and firewalls restricting connections).
  • Execution: An agent performs actions in response to requests from clients or other agents. Agent control loops may run in perpetuity or exit before the end of the agent’s lifetime. Agents may be launched using different mechanisms that are dependent on the application or environment.
These requirements are an extension of actor systems; therefore, our system inherits properties including simplified concurrency, message processing ordering, loose coupling, error isolation, and modularity (Hewitt et al., 1973).
Our implementation focuses on mechanism rather than policy. That is, we discuss how applications can use Academy to achieve certain outcomes without prescribing what agents, or more generally, applications should do. This avoids constraining, or worse, alienating possible use cases and results in a flexible framework suitable for solving many disparate problems. Further, we describe the components within the architecture in terms of abstract interfaces (i.e., without mandating implementation details such as message protocols, state formats, or ordering) to enable further experimentation and optimization, but we still aim to provide implementations that are suitable for most use cases (as demonstrated in the evaluation). Features such as fault tolerance and resilience, resource allocation, and authentication and authorization, while important, are not listed as explicit requirements because applications have varying demands that preclude one-size-fits-all solutions.
3.2. Academy Architecture
@@
Figure 2. Agents and clients in Academy interact via handles to invoke actions asynchronously. Agents implement a behavior, defined by their actions, control loops, and state. Academy decouples the control and data planes through the launcher and exchange components that manage spawning agents and communication, respectively.
[…]
[Academy system architecture.]
As depicted at a high level in Figure 2, an Academy deployment includes one or more agents and zero or more clients. An agent is a process that executes a behavior, where a behavior is defined by a local state, a set of actions, and a set of control loops. Agents are executed remotely using a launcher. Once running, an agent concurrently executes all of its control loops and listens for messages from clients, which can be other agents or programs.
A client interacts with an agent through a handle, a term we borrow from actor frameworks. A handle acts like a reference to the remote agent and translates method calls into action request messages. Each entity (i.e., client or agent) has an associated mailbox that maintains a queue of messages sent to that entity by other entities. Mailboxes are maintained by an exchange such that any client with access to a given exchange can send messages to the mailbox of another agent in the exchange and receive a response through its own mailbox.
3.3. Academy Implementation Details
Academy is implemented as an open-source Python library, available on GitHub and PyPI. [^1] https://github.com/proxystore/academy We target Python for its broad compatibility with scientific workflow codes and libraries, but both the architecture and individual components could be implemented in other languages.
1import time, threading
2from academy.behavior import Behavior, action, loop
@@ 3.3. Academy Implementation Details
4class Example(Behavior):
@@
9 def square(self, value: float) -> float:
@@
13 def count(self, shutdown: threading.Event) -> None:
@@
16 time.sleep(1)
Listing 1: Example agent behavior definition.
3.3.1. Behavior
@@
An agent behavior is implemented as a Python class that inherits from the base Behavior type, as shown in 1. This class-based approach is simple, so existing code can be easily transformed into agents, and extensible through inheritance and polymorphism. Instance attributes maintain the agent’s state, and methods define the actions and control loops.
@@
The @action decorator marks a method as an action, allowing other entities to invoke it remotely. (In the future, we plan to support adding metadata to the @action behavior to aid discovery discussed in Section 3.3.4.) A behavior can invoke actions on itself, as actions are simply Python methods. Methods not decorated as @action are private to the behavior. The @loop decorator marks a method as a control loop. Control loops are executed in separate threads, so a shared threading.Event is passed as an argument to each loop that signals when the agent is shutting down so that control loops can gracefully exit. A control loop can terminate early and the agent will remain running. Commonly, control loops are used to execute a routine on a regular interval, such as to check the state of the environment, or in response to an event. We provide two special control loop decorators, @timer and @event, that simplify behavior implementations for these scenarios.
@@
Two special methods, on_setup() and on_shutdown(), allow behaviors to define callbacks when starting or shutting down, such as to load/store state or initialize/destroy resources. Multiple inheritance of behaviors enables the creation of composite agents.
3.3.2. Agent

### v3

@@ Abstract
Agentic systems, in which diverse agents cooperate to tackle challenging problems, are exploding in popularity in the AI community. However, existing agentic frameworks take a relatively narrow view of agents, apply a centralized model, and target conversational, cloud-native applications (e.g., LLM-based AI chatbots). In contrast, scientific applications require myriad agents be deployed and managed across diverse cyberinfrastructure. Here we introduce Academy, a modular and extensible middleware designed to deploy autonomous agents across the federated research ecosystem, including HPC systems, experimental facilities, and data repositories. To meet the demands of scientific computing, Academy supports asynchronous execution, heterogeneous resources, high-throughput data flows, and dynamic resource availability. It provides abstractions for expressing stateful agents, managing inter-agent coordination, and integrating computation with experimental control. We present microbenchmark results that demonstrate high performance and scalability in HPC environments. To explore the breadth of applications that can be supported by agentic workflow designs, we also present case studies in materials discovery, astronomy, decentralized learning, and information extraction in which agents are deployed across diverse HPC systems.
11footnotetext: Authors contributed equally to this work.
I Introduction
@@
The desire to automate scientific processes has led to advancements in many fields, from artificial intelligence (AI) [1] and computational workflows [2] to research data management [3] and self-driving laboratories (SDL) [4], but humans typically remain responsible for core aspects of the iterative research cycle, including hypothesis generation, experimental design, code development, and data analysis. Automated components stall as humans must trigger experiments, manage workflows, correct errors, and make menial decisions. This friction increases as the scale and ambition of computational science endeavors grow and leads to inefficient use of experimental and observational facilities, data repositories, and high-performance computing (HPC) systems.
@@
Intelligent agents, either individually or composing larger multi-agent systems (MAS), rather than humans, can be the driving entities of discovery [5]. Agents are independent, persistent, stateful, and cooperative—working together to achieve a predefined goal with only intermittent human oversight. The contemporaneous explosion of interest in multi-agent systems is largely a consequence of advancements in reasoning capabilities of the large language models (LLMs) often used to back AI agents [6, 7, 8]. Expressing components of scientific applications as agents—programs that can perform tasks independently or semi-autonomously on behalf of a user or another agent—is powerful. An agent manages its own local state and exposes a well-defined behavior. Agents can perform human roles in iterative scientific processes [9] or encapsulate research cyberinfrastructure (e.g., computational resources and procedures, experimental instruments, data repositories) [10].
Figure 1: Cooperative agents, spanning federated research infrastructure (experimental facilities, computational systems, data storage), can enable agentic workflows that autonomously steer discovery.
However, current agent frameworks (e.g., AutoGen [11]) are not ready to build and deploy agents equipped for scientific applications. They are limited in scope and typically apply a centralized model and target conversational, cloud-native applications (e.g., LLM-based AI chatbots) [12, 11, 13]. Scientific applications, in contrast, typically span geographically (and administratively) separated instruments, supercomputers, robotic facilities, and other resources that collectively constitute federated research infrastructure [14]. Federation creates unique challenges: distributed resources have diverse access protocols, interactions between resources are asynchronous, and varying resource availability requires fault-tolerant and adaptive systems. Existing agentic frameworks are not designed to address these intricacies and workflow management tools designed for federated resources (e.g., Globus Compute [15]) cannot support the long-running and stateful properties of agents natively. New middleware is needed to enable agentic workflows that seamlessly integrate experiment, observation, simulation, analysis, and more, as in Figure 1.
We introduce Academy, a novel actor-based framework for building agentic, scientific workflows, emphasizing modularity, statefulness, and interoperability across federated computing infrastructure. Specifically, this work contributes:
@@
  • Academy, a novel, modular, and extensible middleware for expressing agentic workflows and deploying multi-agent systems across federated resources. Academy addresses unique challenges in scientific applications, such as high data volumes, variable resource availability, and the heterogeneous nature of experimental and computational systems (Section III).
@@
  • Performance analysis of Academy in diverse scenarios yielding insights into the scalability and practical considerations of deploying agentic workflows (Section IV).
@@
  • Case studies demonstrating the utility of agentic workflow design and highlighting improvements in automation, resource utilization, and discovery acceleration (Section V).
@@ II Background
Agents encompass a rapidly expanding front for AI research that can address a breadth of challenges across the computational sciences. We begin with a definition of an agent—inspired by prior work—that is sufficiently generic to encompass the various semantic uses of the term. Then, we describe applications that inform the design of Academy, and distill requirements to support agentic scientific workflows.
@@
An agent is a program that can perform actions independently or semi-autonomously on behalf of a user or another agent. This definition incorporates both the modern notion of intelligent LLM agents [11] and the more traditional definitions of agents [16, 17, 18, 19]. While imprecise, it presents a powerful conceptual model for distributed computing. The agent concept originates from the actor model, a concurrent computing paradigm in which actors encapsulate a local state and communicate through message passing [20]. Agents extend the actor model with the notion of agency—the ability of the agent to engage independently with its environment.
II-A Use Cases
As scientists leverage more advanced computational and experimental resources, and pair them with increasingly capable machine learning models, such as LLMs, the nature of computational science workflows is changing. We highlight applications across four emerging patterns that are guiding our development of agentic middleware.
Steering Applications
Scientists increasingly want to build applications that delegate the direction and composition of scientific campaigns to ML models. Agents simplify the construction of these applications by observing and learning from the results of previous actions and adapting to meet the science goals. For example, Colmena is an active learning library for steering simulation campaigns to discover molecules with desirable properties [21]. In micro-genomics, scientists are using agents to learn the performance of viral identification tools and compose workflows based on sample characteristics [22]. In physical simulations, agentic systems enable on-the-fly learning that dynamically switches between expensive subroutines and learned surrogate models based on the model uncertainty [23].
Decomposing Applications across Facilities
Multi-site workflows are becoming necessary to support complex and heterogeneous scientific applications. These applications need to operate autonomously and manage local state while coordinating workflows across administrative domains. MOFA [24] is a materials science application designed to discover novel metal organic frameworks (MOFs) for carbon capture. It uses a chemical foundation model to generate candidate MOFs and several simulation tools to filter the candidates. The tools are best distributed across multiple computing facilities to meet the heterogeneous resource requirements. Agents can assist at managing the resources allocated at each facility. The Coalition for Epidemic Preparedness Innovations aims to shorten the timeline between disease outbreak and vaccine development to 100 days, which requires coordinating analytics across regions, continuously monitoring outbreaks and trial results, and integrating development and manufacturing into workflows [25]. Agents can facilitate data, information and resource sharing, and automate components of the workflow.
Integrating Instruments into Workflows
As experiments become more data-intensive, scientific applications integrate instruments into computational workflows, or analogously use ML models to control experimental facilities [26]. Embodied agents can integrate instruments into workflows and provide a mechanism to distribute control to the instrument or experimental site. Integral field spectroscopy relies on precise data calibration to study distant spatially complex objects. By deploying agents with the instruments, scientists can track the optical parameters necessary for data-processing and accelerate discovery and steering of the telescope [27]. Self-Driving Laboratories (SDLs) provide an additional tool in chemistry and biology to automatically synthesize materials and measure properties [28]. Distributing agents to SDLs allows on-site control of experiments and integration of these experiments into computational workflows.
Conversational Research Assistants
Scientists are building conversational assistants that help navigate literature, code, documentation, and other data. These systems consist of multiple LLM-based agents that interact (e.g., via dialogue or shared context), with different agents being assigned different roles or capabilities. For instance, Dr. MACS is an astronomy research assistant augmented with retrieval of astronomy literature [29], and ChemGraph is a general purpose chemistry assistant that delegates the construction of chemistry simulation workflows to an LLM [30]. While constructed today with existing agent frameworks, these applications require access to scientific infrastructure to achieve the autonomy and performance required to meet scientific goals.
II-B Requirements to Support Federated Agents
The case studies illuminate benefits of building agentic workflows in science, but also introduce new requirements unmet by existing agent, actor, or workflow systems.
  1. (R1)
    Federated Orchestration. Many of the use-cases require launching and managing agents across different computing systems or scientific instruments.
  2. (R2)
    Configurable Data Plane. Agents may need to frequently communicate at the pace of simulation time-steps on high-performance interconnects or large volumes between facilities and instruments for long running campaigns. The communication mechanism that agents use needs to be configurable to adapt to the specific application.
  3. (R3)
    Temporally Decoupled Messaging. Research infrastructure has varying availability, typically with much lower uptimes than cloud infrastructure [31]. Communication between agents must cope with facilities being temporarily unavailable.
  4. (R4)
    Agent Authentication and Permissions. Agents with the capability to use research infrastructure risk exposing the infrastructure to untrusted users. Researchers must be able to securely delegate (and revoke) the permissions to use tools and infrastructure to agents.
  5. (R5)
    Resilient State Management. State is a key feature across use-cases allowing these agentic systems to learn and adapt. While state is a necessary feature of the actor model, it is ill supported across workflow management tools that are popular among scientific applications.
The adoption of specific frameworks is often also influenced by usability requirements such as intuitive and simple representations, local testing and debugging capabilities, etc.
III AcademyDesign
In the design of Academy, we aim to address the following high-level challenges: How to represent, in code, the declaration of and interaction between agents? How to deploy agents across federated infrastructure? How to achieve performance across heterogeneous systems and networks? Academy is an open-source Python library, available on GitHub [^1] https://github.com/academy-agents and PyPI. We target Python for its broad compatibility with scientific workflow codes and libraries, but both the architecture and individual components could be implemented in other languages.
III-A Academy Architecture
@@
Figure 2: Agents and users in Academy interact via handles to invoke actions asynchronously. Agents implement a behavior, defined by their actions, control loops, and state. Academy decouples the control and data planes through the executor and exchange components that manage spawning agents and communication, respectively.
As depicted at a high level in Figure 2, an Academy deployment includes one or more agents and zero or more users. An agent is a process with a local state, a set of actions, and a set of control loops. Agents are executed remotely using an executor. Once running, an agent concurrently executes all of its control loops and listens for messages from clients, which can be other agents or users.
A client interacts with an agent through a handle, a term we borrow from actor frameworks. A handle acts like a reference to the remote agent and translates method calls into action request messages. Each entity (i.e., user or agent) has an associated mailbox that maintains a queue of messages sent to that entity by other entities. Mailboxes are maintained by an exchange such that any client with access to a given exchange can send messages to the mailbox of another agent in the exchange and receive a response through its own mailbox.
III-B Academy Interaction
We first describe how scientists program agents and interact with the Academy ecosystem.
III-B1 Agent Representation
1from asyncio import Event, sleep
2from academy.agent import Agent, action, loop
@@ III-B1 Agent Representation
4class Example(Agent):
@@
9 async def square(self, value: float) -> float:
@@
13 async def count(self, shutdown: Event) -> None:
@@
16 await sleep(1)
Listing 1: Example agent definition, showing initialization that sets an internal count variable, an action that squares a supplied value, and a loop that increments the internal count variable once a second.
@@
An agent is implemented as a Python class that inherits from the base Agent type: see 1. This class-based approach is simple, so existing code can be easily transformed into agents, and extensible through inheritance and polymorphism. Instance attributes maintain the agent’s state, and methods define the actions and control loops.
@@
The @action decorator marks a method as an action, allowing other entities to invoke it remotely. An agent can invoke actions on itself, as actions are simply Python methods. Methods not decorated as @action are private to the agent. The @loop decorator marks methods as control loops. Control loops extend the actor model to enable the programming of autonomous behavior. Control loops are executed concurrently in an event loop; a shared Event is passed to each loop to signal agent shutdown, so that control loops can exit gracefully. A control loop can terminate early and the agent will remain running. Commonly, control loops are used to execute a routine on a regular interval, such as to check the state of the environment, or in response to an event. We provide special loop decorators, such as @timer and @event, that simplify agent implementations for common use cases.
@@
Two special methods, agent_on_startup() and agent_on_shutdown(), allow agents to define callbacks when starting or shutting down, such as to load/store state or initialize/destroy resources. Multiple inheritance of agents enables the creation of composite agents.
III-B2 Agent Invocation
A handle is a client interface to a remote agent used to invoke actions, ping, and shutdown the agent. Each handle acts as a reference to that agent, translating each method call into a request message that is sent via the exchange and asynchronously waiting on the response message and returning the result. The handle decides which mailbox to send from and listen to based on the context where the handle is used; a handle can be passed between users and agents and automatically attaches to the mailbox of the respective client. This ensures that there is only one listening task per mailbox, and one mailbox per client (i.e., agent or user).
III-B3 Manager Class
A Manager combines an exchange and one or more executors to provide a single interface for launching, using, and managing agents. This reduces boilerplate code, improves communication efficiency, and ensures stateful resources and tasks are appropriately cleaned up. An end-to-end example is provided in 2.
III-C Agent Management
III-C1 Agent Runtime